40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 13 01:18:00 2014, Jenne, Update, LSC, Composite Error Signal for ARms (8) 7x
    Reply  Fri Mar 14 12:18:55 2014, Koji, Update, LSC, Composite Error Signal for ARms (9) asymptotic_cooling.pdf
Message ID: 9724     Entry time: Thu Mar 13 01:18:00 2014     Reply to this: 9728
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Composite Error Signal for ARms (8) 

[Jenne, EricQ]

As Koji suggested in his email this afternoon, we looked at how much actuator range is required for various forms of arm locking:  (1) "regular" PDH lock aquisition, (2) ALS lock acquisition, (3) CESAR cooling.

To start, I looked at the spectra and time series data of the control signal (XARM_OUT) for several locking situations.  Happily, when the arm is at the half fringe, where we expect the 1/sqrt(TRX) signal to be the most sensitive (versus the same signal at other arm powers), we see that it is in fact more quiet than even the PDH signal.  Of course, we can't ever use this signal once the arm is at resonance, so we haven't discovered anything new here.

XARM_OUT_VariousErrorSignals_ungarb.pdf

EricQ then made some violin plots with the time series data from these situations, and we determined that a limit of ~400 counts encompasses most of the steady-state peak-to-peak output from locking on the PDH signal.

xarmOutViolinPlot.pdfxarmOutViolinSub.pdf

[ericq: What's being plotted here are "kernel density estimates" of the time series data of XARM_OUT when locked on these signals. The extent of the line goes to the furthest outlier, while the dashed and dotted lines indicate the median and quartiles, respectively]

I tried acquiring ALS and transitioning to final PDH signals with different limiters set in the Xarm servo.  I discovered that it's not too hard to do with a limit of 400 counts, but that below ~350 counts, I can't keep the ALS locked for long enough to find the IR resonance.  Here's a plot of acquiring ALS lock, scanning for the resonance, and then using CESAR to transition to PDH, with the limit of 400 counts in place, and then a lockloss at the end.  Even though I'm hitting the rails pretty consistently, until I transition to the more quiet signals, I don't ever lose lock (until, at the end, I started pushing other buttons...).

LimiterAt400cts.pdf

After that, I tried acquiring lock using our "regular" PDH method, and found that it wasn't too hard to capture lock with a limit of 400, but with limits below that I can't hold the lock through the boosts turning on.

noLimitPDHAcq.pdfwithLimitPDHAcq.pdf

Finally, I took spectra of the XARM_OUT control signal while locked using ALS only, but with different limiter values. Interestingly, I see much higher noise between 30-300 Hz with the limiter engaged, but the high frequency noise goes down.  Since the high frequency is dominating the RMS, we see that the RMS value is actually decreasing a bit (although not much).

XARM_OUT_VariousLimits_ungarb.pdf

We have not made any changes to the LSC model, so there is still no smoothing between the ALS and IR signals.  That is still on the to-do list.  I started modifying things to be compatible with CARM rather than a single arm, but that's more of a daytime-y task, so that version of the c1lsc model is saved under a different name, and the model that is currently compiled and running is reverted as the "c1lsc.mdl" file.

ELOG V3.1.3-